# Sharp asymptotics for $N$-point correlation functions of coalescing heavy-tailed random walks

### v1

@@ 4.1 $N$-point correlation function
For ${\bf x}\in{\mathbb{Z}}^{N}$, recall the $N$-point function $\rho_{N}({\bf x},t)$ from (1.8) and recall the density function $\rho_{1}(t)$. For the CRW $(\xi_{t})_{t\geq 0}$ starting from every site on ${\mathbb{Z}}$ and for any $\textbf{w}=(w_{1},\ldots,w_{n})\in{\mathbb{Z}}^{n}$, with a bit abuse of notation, we define $\xi_{t}(\textbf{w}):=\prod_{i=1}^{n}\xi(w_{i})\in\{0,1\}$. Then, we can rewrite that $\rho_{N}({\bf x},t)={\mathbb{E}}[\xi_{t}({\bf x})]$. In this subsection, we denote $p_{\rm NC}(t):={\mathbb{P}}(\hat{\tau}_{\bf x}>t)$.
@@
${\mathbb{E}}[\xi_{t}({\bf x})]=\rho_{1}(t-s)^{N}\hskip 1.0ptp_{\rm NC}(s)\big{(}1+{\cal R}_{1}+{\cal R}_{2}+{\cal R}_{3}\big{)},$ (4.1)
[…]
@@
[…]
Proof Firstly, we trace back to time $t-s$ for the ancestors of the $N$ particles at ${\bf x}$ at time $t$. For ${\bf y}\in{\mathbb{Z}}^{N}$, we say that $({\bf y},t-s)$ connects to $({\bf x},t)$ if $N$ coalescing random walks in the system move from ${\bf y}$ at time $t-s$ to ${\bf x}$ at $t$. By the natural coupling between the coalescing random walks and independent random walks, and by the time-reversal argument,
@@
By the inclusion-exclusion principle, $R_{1}\leq 0$ and $|R_{1}|$ is bounded above by the probability of the following event: there exist ${\bf y}\neq{\bf z}\in{\mathbb{Z}}^{N}$ such that (a) $\xi_{t-s}({\bf y},{\bf z})=1$ (i.e., there are particles at all sites of $\{{\bf y},{\bf z}\}$ at time $t-s$), and (b) both $({\bf y},t-s)$ and $({\bf z},t-s)$ connect to $({\bf x},t)$. W.l.o.g., we rearrange the order of coordinates, and assume that the first $J$ coordinates of ${\bf y}$ and ${\bf z}$ differ while the remaining $N-J$ coordinates are the same. Since then the two random walks starting from $y_{i}$ and $z_{i}$ at $t-s$ must coalesce before time $t$, assume additionally that their coalescing times are ordered $t-s<u_{J}<\ldots<u_{1}<t$. To obtain an upper bound for the probability of (b), we consider the time-reversed random walks and set $\hat{u}_{i}:=t-u_{i}$. For the reversed time from 0 to $\hat{u}_{1}$, $N$ random walks $(\widehat{\mathbf{X}}_{t})_{t\geq 0}$ starting from ${\bf x}$ does not collide. For $1\leq i\leq J$, from $\hat{u}_{i}$ to $\hat{u}_{i+1}$ (with $\hat{u}_{J+1}:=s$), we compare with the forward coalescing system. Given $\widehat{X}^{i}_{\hat{u}_{i}}$, there is a jump to it at $\hat{u}_{i}$ from another walk with $\widehat{Y}^{i}_{\hat{u}_{i}}=\widehat{X}^{i}_{\hat{u}_{i}}+y_{i}$, and the pair does not collide by time $\hat{u}_{i+1}$. This yields an upper bound
∫⋯∫$\idotsint\displaylimits_{0<\hat{u}_{1}<\hat{u}_{2}<\ldots<\hat{u}_{J}<s}\sum_{{\bf x},y_{i}}{\mathbb{P}}_{\bf x}(\hat{\tau}_{{\bf x}}>\hat{u}_{1})p(y_{1}){\mathbb{P}}_{y_{1}}(\tau_{y_{1}}>\hat{u}_{1})\ldots p(y_{J}){\mathbb{P}}_{y_{J}}(\tau_{y_{J}}>\hat{u}_{J})\mathrm{d}\hat{u}_{1}\ldots\mathrm{d}\hat{u}_{J}.∫⋯$
By Theorem 1.4 Lemma 2.9, this integral is bounded above by
$C\hskip 1.0pt{\mathbb{P}}_{\bf x}(\hat{\tau}_{{\bf x}}>s)\big{(}s/l(s)\big{)}^{J}\leq C_{\bf x}\hskip 1.0ptp_{\rm NC}(s)\hskip 1.0pt\big{(}s/l(s)\big{)}^{J}.$ (4.2)
@@
For (a), since there are exactly $N+J$ particle at $\{{\bf y},{\bf z}\}$ at $t-s$, by the negative correlation of CRW (see Theorem A.1),
${\mathbb{E}}[\xi({\bf y},{\bf z})]\leq\rho_{1}(t-s)^{N+J}.$ (4.3)
Note that $1\leq J\leq N$ and $\big{(}s\hskip 1.0pt\rho_{1}(t-s)/l(s)\big{)}^{J}$ is either bounded by $s\hskip 1.0pt\rho_{1}(t-s)/l(s)$ or by $\big{(}s\hskip 1.0pt\rho_{1}(t-s)/l(s)\big{)}^{N}$. Combining (4.2) and (4.3), we have
$|R_{1}|\leq C_{\bf x}\sum_{J=1}^{N}p_{\rm NC}(s)\hskip 1.0pt\Big{(}\frac{s}{l(s)}\Big{)}^{J}\rho_{1}(t-s)^{N+J}=\rho_{1}(t-s)^{N}\hskip 1.0ptp_{\rm NC}(s)O\Big{(}\sum_{J=1}^{N}\Big{(}\frac{s\hskip 1.0pt\rho_{1}(t-s)}{l(s)}\Big{)}^{J}\Big{)},$
which gives rise to ${\cal R}_{1}$ in (4.1).
@@
By the negative correlation ${\mathbb{E}}[\xi_{t-s}]\leq\rho_{1}(t-s)^{N}$, the asymptotic independence Proposition 3.3 and Remark 3.4 then imply that there exists $\delta>0$ such that
$|R_{2}|\leq\frac{C_{\bf x}\hskip 1.0pt\rho_{1}(t-s)^{N}}{l(s)^{\upsilon_{N}}(\ell(s)^{1-\varepsilon}\wedge t^{\delta})}\leq\rho_{1}(t-s)^{N}\hskip 1.0ptp_{\rm NC}(s)\frac{C_{\bf x}}{(\ell(s)^{1-\varepsilon}\wedge t^{\delta})},$ (4.4)
where we used Theorem 1.4 for $p_{\rm NC}$ in the last inequality. This leads to ${\cal R}_{2}$ in (4.1).
@@
For the density function, Theorem B.1 gives a priori bound $\frac{c}{t}\leq\rho_{1}(t)\leq\frac{C\hskip 1.0ptl(t)}{t}$ for some constants $c,C>0$. With the notation that $\xi(0,x)=\xi(0)\xi(x)$, the ODE (1.13) obtained via the generator calculation can be rewritten as
$-\frac{\mathrm{d}}{\mathrm{d}t}\rho_{1}(t)=\sum_{x}p(x){\mathbb{E}}[\xi_{t}(0,x)].$ (4.5)
With these two ingredients, we show Theorem 1.2 in three steps.
The first step is to improve the lower bound: there exists $c>0$ such that
$\rho_{1}(t)\geq\frac{c\hskip 1.0ptl(t)}{t}.$ (4.6)
@@
We only need to consider the case $l(\infty)=\infty$ when $\alpha=1$, since otherwise the result is trivial. As in Lemma 4.1 ${\cal R}_{1}\leq 0$ and ${\cal R}_{2}\leq C$ for some $C>0$, we have that for $s<t$,
$-\frac{\mathrm{d}}{\mathrm{d}t}\rho_{1}(t)\leq C\rho_{1}(t-s)^{2}\sum_{x}p(x){\mathbb{P}}(\hat{\tau}_{x}>s)\hskip 1.0pt\Big{(}1+\frac{1}{B_{s}\hskip 1.0pt\rho_{1}(t-s)}\Big{)}.$ (4.7)
@@
Recall that $B_{t}\in{\cal B}$, there exists $K$ such that $\ell(t)^{K}\geq l(t)$. Let $\theta>1-1/K$. For each $t$, define $s=s(t)$ from the following equation:
$s\hskip 1.0pt\rho_{1}(t-s)=l(s)^{\theta}.$ (4.8)
@@
By a priori upper bound of $\rho_{1}$, for $s=t^{\delta}$ with $\delta<1$, the l.h.s. of (4.8) is bounded above by $t^{-(1-\delta)+\varepsilon}\ll l(s)^{\theta}$. We thus conclude $t^{\delta}<s<t$. Moreover, for any $\varepsilon>0$,
$t/s=t\hskip 1.0pt\rho_{1}(t-s)\hskip 1.0ptl(s)^{-\theta}\leq l(t)\hskip 1.0ptl(s)^{-\theta}\leq l(t)^{1-\theta}(t/s)^{\varepsilon}.$ (4.9)
It also follows that $-\rho^{\prime}_{1}(t)\leq C\rho_{1}(t-s)^{2}/l(s)$ by (4.7) and Lemma 2.9 and the fact that
$\frac{1}{B_{s}\hskip 1.0pt\rho_{1}(t-s)}=\frac{l(s)^{1-\theta}}{\ell(s)}\leq\ell(s)^{-1+K(1-\theta)}\underset{t\to 0}{\longrightarrow}0.$ (4.10)
@@
To simplify the ODE, we have to replace $\rho_{1}(t-s)^{2}$ and $l(s)$. By Lemma 2.10 and (4.9), we have $1/l(s)-1/l(t)\leq Cl(t)^{-1+\varepsilon}\ell(t)^{-1}$ for any $\varepsilon>0$. Furthermore, we use (B.8) with a change of $t$ to $t-s$, and take $L=2/\rho_{1}(t-s)$ therein. It follows that $\frac{Cs}{B_{s}^{2}l(s)\rho_{1}(t-s)}\leq\frac{C}{\ell(s)B_{s}\rho_{1}(t-s)}<1$, Thus (B.8) implies that $\rho_{1}(t-s)\leq C\rho_{1}(t)$. Therefore, by bounding the r.h.s. of (4.7) we obtain
$-\frac{\mathrm{d}}{\mathrm{d}t}\rho_{1}(t)\leq C\rho_{1}(t)^{2}/l(t),$
which yields the lower bound (4.6).
The second step is to drive the sharp bounds in Theorem 1.2. Recall Lemma 4.1 with $N=2$ and error terms ${\cal R}_{1},{\cal R}_{2},{\cal R}_{3}$. By (4.5), we have
$-\frac{\mathrm{d}}{\mathrm{d}t}\rho_{1}(t)=\rho_{1}(t-s)^{2}\sum_{x}p(x){\mathbb{P}}(\hat{\tau}_{x}>s)\hskip 1.0pt\big{(}1+{\cal R}_{1}+{\cal R}_{2}+{\cal R}_{3}\big{)}.$ (4.11)
@@
Since the error terms have different expressions for $\alpha=1$ and $\alpha<1$, for convenience below we focus on $\alpha=1$ and $l(\infty)=\infty$, namely the recurrent case. But it is easy to check that all arguments, with slight modifications, work for $\alpha<1$. Again take $s$ as in (4.8) with $\theta\in(1-1/K,1)$. It follows that
${\cal R}_{1}=O\Big{(}\frac{1}{l(t)^{1-\theta}}\Big{)},\quad{\cal R}_{2}=O\Big{(}\frac{1}{\ell(t)^{1-\varepsilon}}\Big{)},\quad\mbox{and}\quad{\cal R}_{3}=O\Big{(}\frac{l(s)^{1-\theta}}{\ell(s)}\Big{)}.$ (4.12)

### v2

@@ 4.1 $N$-point correlation function
For ${\bf x}\in{\mathbb{Z}}^{N}$, recall the $N$-point function $\rho_{N}({\bf x},t)$ from (12) and recall the density function $\rho_{1}(t)$. For the CRW $(\xi_{t})_{t\geq 0}$ starting from every site on $\mathbb{Z}$} and for any $\textbf{w}=(w_{1},\ldots,w_{n})\in{\mathbb{Z}}^{n}$, with a bit abuse of notation, we define $\xi_{t}(\textbf{w}):=\prod_{i=1}^{n}\xi_{t}(w_{i})\in\{0,1\}$. Then, we can rewrite that $\rho_{N}({\bf x},t)={\mathbb{E}}[\xi_{t}({\bf x})]$. In this subsection, we denote $p_{\rm NC}(t):={\mathbb{P}}(\hat{\tau}_{\bf x}>t)$.
@@
${\mathbb{E}}[\xi_{t}({\bf x})]=\rho_{1}(t-s)^{N}\hskip 1.0ptp_{\rm NC}(s)\big(1+{\cal R}_{1}+{\cal R}_{2}+{\cal R}_{3}\big),$ (91)
where $0\geq{\cal R}_{1}=O\Big(\sum\limits_{J=1}^{N}\big(\frac{s\hskip 1.0pt\rho_{1}(t-s)}{l(s)}\big)^{J}\Big)$, ${\cal R}_{2}=O\Big(\frac{1}{\ell(t)^{1-\varepsilon}\wedge t^{\delta}}\Big)$ and $0\leq{\cal R}_{3}=O\Big(\sum\limits_{J=1}^{N-1}\big(\frac{1}{B_{s}\hskip 1.0pt\rho_{1}(t-s)}\big)^{J}\Big)$.
@@
Later when we use Lemma 4.1, we choose $s$ such that $\frac{s\hskip 1.0pt\rho_{1}(t-s)}{l(s)},\frac{(t/s)^{\varepsilon}}{B_{t}\hskip 1.0pt\rho_{1}(t-s)}\ll 1$, and thus ${\cal R}_{1}$ and ${\cal R}_{3}$ can be written in much tighter forms.
Proof Firstly, we trace back to time tst-s for the ancestors of the $N$ particles at $\bf x$} at time $t$. For ${\bf y}\in{\mathbb{Z}}^{N}$, we say that $({\bf y},t-s)$ connects to $({\bf x},t)$ if $N$ coalescing random walks in the system move from $\bf y$} at time tst-s to $\bf x$} at $t$. By the natural coupling between the coalescing random walks and independent random walks, and by the time-reversal argument,
@@
Indeed, if we call $A({\bf y})$ the event that in the coalescing system there are particles at $\bf y$} at time tst-s which connects to $\bf x$} at time $t$, then in the equation above, the l.h.s. equal ${\mathbb{P}}(\cup_{\bf y}A({\bf y}))$ and the sum in the r.h.s. equals $\sum_{\bf y}{\mathbb{P}}(A({\bf y}))$. Thus by the inclusion-exclusion principle, $R_{1}\leq 0$ and $|R_{1}|$ is bounded above by the probability of the following event: there exist ${\bf y}\neq{\bf z}\in{\mathbb{Z}}^{N}$ such that (a) $\xi_{t-s}({\bf y},{\bf z})=1$ (i.e., there are particles at all sites of $\{{\bf y},{\bf z}\}$ at time tst-s), and (b) both $({\bf y},t-s)$ and $({\bf z},t-s)$ connect to $({\bf x},t)$. Suppose that there are $J$ different coordinates of $\bf y$} and $\bf z$} and the remaining NJN-J coordinates are the same. W.l.o.g., we rearrange the order of coordinates, and assume that the first $J$ coordinates of $\bf y$} and $\bf z$} differ. Since then the two random walks starting from $y_{i}$ and $z_{i}$ at tst-s must coalesce before time $t$ for $1\leq i\leq J$, assume additionally that their coalescing times are ordered $t-s<u_{J}<\ldots<u_{1}<t$. In words, (a) is the event that in the coalescing system there are $N+J$ particles occupying the sites $\{{\bf y},{\bf z}\}=\{y_{1},\ldots,y_{N},z_{1},\ldots,z_{J}\}$ at time tst-s, and (b) is the event that the pair of particles at site $y_{i}$ and $z_{i}$ move to $x_{i}$ at time $t$ for $1\leq i\leq J$ and the particle at $y_{j}$ moves to $x_{j}$ at time $t$ for $J<j\leq N$. $|R_{1}|$ is then bounded from above by summing the probabilities of (a) intersects (b) for all $\{{\bf y},{\bf z}\}$ over ${\mathbb{Z}}^{N+J}$ and then for $J$ from $1$ to $N$.
@@
To obtain an upper bound for the probability of (a), we recall that there are exactly $N+J$ particle at $\{{\bf y},{\bf z}\}$ at tst-s, and by the negative correlation of CRW (see Theorem A.1),
${\mathbb{E}}[\xi({\bf y},{\bf z})]\leq\rho_{1}(t-s)^{N+J}.$ (92)
We then take (b) into account. It implies that the following event on the time-reversed (branching) random walks occurs. Set $\hat{u}_{i}:=t-u_{i}$. For the reversed time from 0 to $\hat{u}_{1}$, $N$ random walks $(\widehat{\mathbf{X}}_{t})_{t\geq 0}$ starting from $\bf x$} does not collide. For $1\leq i\leq J$, from $\hat{u}_{i}$ to $\hat{u}_{i+1}$ (with $\hat{u}_{J+1}:=s$), we compare with the forward coalescing system. Given $\widehat{X}^{i}_{\hat{u}_{i}}$, there is a binary branching at time $\hat{u}_{i}$ due to the coalescence in the forward coalescing system, where one descendant stays and the other jumps to $\widehat{Y}^{i}_{\hat{u}_{i}}=\widehat{X}^{i}_{\hat{u}_{i}}+w_{i}$ with probability $p(w_{i})$, and the pair does not collide at least by time $\hat{u}_{i+1}$. We note that the branching rate is $2$ as in the forward system either particle in $i$-th pair can jump and coalesce with the other. Since (92) gives a uniform bound, for the upper bound of $|R_{1}|$, the sum of $\{yb,{\bf z}\}$ is applied to (b) only, and it is equivalent to that there is no restriction for the time-reversed random walks at reversed time $s$. This yields an upper bound of $|R_{1}|$ as
$\sum_{J=1}^{N}\rho_{1}(t-s)^{N+J}\hskip-15.0pt\idotsint\displaylimits_{0<\hat{u}_{1}<\hat{u}_{2}<\ldots<\hat{u}_{J}<s}\hskip-5.0ptC_{\bf x}\hskip-5.0pt\sum_{w_{1},\ldots,w_{J}}\hskip-5.0pt{\mathbb{P}}_{\bf x}(\hat{\tau}_{{\bf x}}>\hat{u}_{1})p(y_{1}){\mathbb{P}}_{y_{1}}(\tau_{y_{1}}>\hat{u}_{1})\ldots p(y_{J}){\mathbb{P}}_{y_{J}}(\tau_{y_{J}}>\hat{u}_{J})\mathrm{d}\hat{u}_{1}\ldots\mathrm{d}\hat{u}_{J}.$
By Theorem 1.2 and Lemma 2.9, the multiple integral in the above display is bounded from above by
$C_{\bf x}\hskip 1.0pt{\mathbb{P}}_{\bf x}(\hat{\tau}_{{\bf x}}>s)\big(s/l(s)\big)^{J}=C_{\bf x}\hskip 1.0ptp_{\rm NC}(s)\hskip 1.0pt\big(s/l(s)\big)^{J}.$ (93)
Therefore, we have
$|R_{1}|\leq C_{\bf x}\sum_{J=1}^{N}\rho_{1}(t-s)^{N+J}p_{\rm NC}(s)\hskip 1.0pt\Big(\frac{s}{l(s)}\Big)^{J}=\rho_{1}(t-s)^{N}\hskip 1.0ptp_{\rm NC}(s)O\Big(\sum_{J=1}^{N}\Big(\frac{s\hskip 1.0pt\rho_{1}(t-s)}{l(s)}\Big)^{J}\Big),$
which gives rise to ${\cal R}_{1}$ in (91).
@@
By the negative correlation ${\mathbb{E}}[\xi_{t-s}({\bf y})]\leq\rho_{1}(t-s)^{N}$, the asymptotic independence Proposition 3.3 and Remark 3.4 then imply that there exists $\delta>0$ such that
$|R_{2}|\leq\frac{C_{\bf x}\hskip 1.0pt\rho_{1}(t-s)^{N}}{l(s)^{\upsilon_{N}}(\ell(s)^{1-\varepsilon}\wedge t^{\delta})}\leq\rho_{1}(t-s)^{N}\hskip 1.0ptp_{\rm NC}(s)\frac{C_{\bf x}}{(\ell(s)^{1-\varepsilon}\wedge t^{\delta})},$ (94)
where we used Theorem 1.2 for $p_{\rm NC}$ in the last inequality. This leads to ${\cal R}_{2}$ in (91).
@@
To give precise asymptotics of the density, we need a priori rough bounds of $\rho_{1}$ that give the correct order, where the upper bound can be proved using a modification of the arguments of BG (80).
Proposition 4.3
$\frac{c\hskip 1.0ptl(t)}{t}\leq\rho_{1}(t)\leq\frac{C\hskip 1.0ptl(t)}{t}.$ (95)
Proof The proof proceeds in three steps, establishing successively a crude lower bound $\rho_{1}(t)\geq c/t$, an upper bound $\rho_{1}(t)\leq C\hskip 1.0ptl(t)/t$, and finally an improved lower bound $\rho_{1}(t)\geq c\hskip 1.0ptl(t)/t$.
Step 1. Recall the notation that $\xi(0,x)=\xi(0)\xi(x)$. By the generator calculation, using the negative correlation ${\mathbb{E}}[\xi(0,x)]\leq{\mathbb{E}}[\xi(0)]{\mathbb{E}}[\xi(x)]$ (see Theorem A.1), and by the translation invariance ${\mathbb{E}}[\xi(x)]=\rho_{1}(t)$,
$\frac{\mathrm{d}}{\mathrm{d}t}\rho_{1}(t)=\sum_{x\in{\mathbb{Z}}}p(-x){\mathbb{E}}[\xi_{t}(0)\xi_{t}(x)]\leq\sum_{x\in{\mathbb{Z}}}p(-x)\rho_{1}(t)^{2}=\rho_{1}(t)^{2},$ (96)
which yields a crude bound
$\rho_{1}(t)\geq\frac{c}{t}.$
Step 2. The idea of proving the upper bound is credited to BG (80). We thus only provide a proof sketch here. The idea is that given the density $\rho_{1}(t)$, starting from the interval of length $C/\rho_{1}(t)$, there is a probability bounded uniformly from below that all other particles would coalesce with one particle over time $s$, which results to a lower bound for the density decay. For any $L\in{\mathbb{N}}$, Denote $\Lambda:=[0,L]\cap{\mathbb{Z}}$ and $x+\Lambda:=\{x+y:y\in\Lambda\}$. Let $\xi^{A}_{t}$ be the coalescing random walks starting from set $A$. Identifying $\xi_{t}(B)$ as the set $\{x\in B:\xi_{t}(x)=1\}$ for any set $B\subseteq{\mathbb{Z}}$, we have that for any $s,t\geq 0$,
[…]
where in the inequality we applied the Markov property at time $t$ for ancestors of the particles in $\Lambda$ at time $t+s$, and in the first equality we used translation invariance. We note that all terms in the equation above are finite as ${\mathbb{E}}\big[|\xi^{\xi_{t}(\Lambda)}_{s}({\mathbb{Z}})|\big]\leq{\mathbb{E}}\big[|\xi_{t}(\Lambda)|\big]\leq|\Lambda|$. Since ${\mathbb{E}}[|\xi_{t}(\Lambda)|]=L\rho_{1}(t)$, we derive
[…]
The right-hand side of (97) is the expected number of times that the particles $\xi_{t}(\Lambda)$ coalesce through time $s$. This is lower bounded if we fix a particle in $\xi_{t}(\Lambda)$, and consider the number of coalescence between this particle and others only, which is further bounded from below by
${\mathbb{E}}\big[\xi_{t}(\Lambda)-1\big]\cdot\min_{|x|\leq L}{\mathbb{P}}(\tau_{x}\leq 2s).$
Recall that $\Delta(x,s)={\mathbb{P}}(X^{0}_{s}=0)-{\mathbb{P}}(X^{x}_{s}=0)$. Using the last-exit decomposition (48),
${\mathbb{P}}(\tau_{x}\leq s)=1-\Delta(x,s)-\int_{0}^{s}\Delta(x,u)j(s-u)\mathrm{d}u.$
Since $|\Delta(x,u)|\leq\frac{c|x|}{B_{u}^{2}}\leq\frac{cL}{B_{u}^{2}}$ and $j(u)\sim 1/l(u)$, it follows $|\int_{0}^{s}\Delta(x,u)j(s-u)\mathrm{d}u|\leq\frac{csL}{B_{s}^{2}l(s)}$. Inserting all these bounds into (97), we get
$L\rho_{1}(t)-L\rho_{1}(t+s)\geq\big(L\rho_{1}(t)-1\big)\Big(1-\frac{csL}{B_{s}^{2}l(s)}\Big),$ (98)
$L\rho_{1}(2t)\leq P^{\kappa}L\rho_{1}(t)+(1-P)^{-1}.$
$\rho_{1}(2t)\leq\frac{\rho_{1}(t)}{\ell(t)}+\frac{2l(t)}{t}.$
Then, for all $t$ such that $\ell(t)>10$,
$\rho_{1}(2t)-\frac{5\hskip 1.0ptl(2t)}{2t}\leq\frac{1}{10}\Big(\rho_{1}(t)-\frac{5\hskip 1.0ptl(t)}{t}\Big),$
where we used $l(2t)\geq l(t)$ for the increasing function $l$. The upper bound in (95) follows easily from the last inequality.
Step 3. To improve the lower bound to
$\rho_{1}(t)\geq\frac{c\hskip 1.0ptl(t)}{t},$ (99)
@@
we only need to consider the recurrent case, namely when $\alpha=1$ and $l(\infty)=\infty$, since otherwise the result is trivial. As in Lemma 4.1 ${\cal R}_{1}\leq 0$ and ${\cal R}_{2}\leq C$ for some $C>0$, we have that for $s<t$,
$\frac{\mathrm{d}}{\mathrm{d}t}\rho_{1}(t)\leq C\rho_{1}(t-s)^{2}\sum_{x}p(x){\mathbb{P}}(\hat{\tau}_{x}>s)\hskip 1.0pt\Big(1+\frac{1}{B_{s}\hskip 1.0pt\rho_{1}(t-s)}\Big).$ (100)
@@
Recall that $B_{t}\in{\cal B}$, there exists $K$ such that $\ell(t)^{K}\geq l(t)$. Let $1>\theta>1-1/K$. For each $t$, define $s=s(t)$ from the following equation:
$s\hskip 1.0pt\rho_{1}(t-s)=l(s)^{\theta}.$ (101)
It follows that $\rho^{\prime}_{1}(t)\leq C\rho_{1}(t-s)^{2}/l(s)$ by (100) and Lemma 2.9 and the fact that
$\frac{1}{B_{s}\hskip 1.0pt\rho_{1}(t-s)}=\frac{l(s)^{1-\theta}}{\ell(s)}\leq\ell(s)^{-1+K(1-\theta)}\underset{t\to 0}{\longrightarrow}0.$ (102)
@@
By the upper bound of $\rho_{1}$, for $s=t^{\delta}$ with $\delta<1$, the l.h.s. of (101) is bounded above by $t^{-(1-\delta)+\varepsilon}\ll l(s)^{\theta}$. We thus conclude $t^{\delta}<s<t$. Moreover, for any $\varepsilon>0$, by the rough upper bound in (95),
$t/s=t\hskip 1.0pt\rho_{1}(t-s)\hskip 1.0ptl(s)^{-\theta}\leq C\hskip 1.0ptl(t)\hskip 1.0ptl(s)^{-\theta}\leq C\hskip 1.0ptl(t)^{1-\theta}(t/s)^{\varepsilon}.$ (103)
To simplify the ODE, we have to replace $\rho_{1}(t-s)^{2}$ and $l(s)$. By Lemma 2.10 and (103), for any $\varepsilon>0$,
$1/l(s)-1/l(t)\leq Cl(t)^{-1+\varepsilon}\ell(t)^{-1}.$ (104)
@@
Furthermore, we use (98) with a change of $t$ to tst-s, and take $L=2/\rho_{1}(t-s)$ therein. It follows that $\frac{Cs}{B_{s}^{2}l(s)\rho_{1}(t-s)}\leq\frac{C}{\ell(s)B_{s}\rho_{1}(t-s)}<1$, Thus (98) implies that $\rho_{1}(t-s)\leq C\rho_{1}(t)$. Therefore, by bounding the r.h.s. of (100) we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\rho_{1}(t)\leq C\rho_{1}(t)^{2}/l(t),$
which yields the lower bound (99).
Proof of Theorem 1.1 We can now derive the sharp bounds in Theorem 1.1. Recall that for each $t$ we define $s=s(t)$ through (101), that is, for $\theta\in(1-1/K,1)$,
$s\hskip 1.0pt\rho_{1}(t-s)=l(s)^{\theta}.$ (105)
On one hand, the monotonicity of $l$ and the fact that $t>s$ imply
$s\hskip 1.0pt\rho_{1}(t-s)\leq l(t)^{\theta}.$
On the other hand, take $\tilde{s}=t/l(\theta)^{\delta}$ with $0<\delta<1-\theta$. Using the rough bound $\rho_{1}(t)\asymp l(t)/t$ from Proposition 4.3, we obtain
$\tilde{s}\hskip 1.0pt\rho_{1}(t-\tilde{s})\geq t\hskip 1.0ptl(t)^{-\delta}c\hskip 1.0ptl(t)/t=c\hskip 1.0ptl(t)^{1-\delta},$
and the r.h.s. is larger than $l(t)^{\theta}$ for $t$ sufficient large. Hence, by monotonicity of $u\mapsto u\rho_{1}(t-u)$, we conclude that
$s(t)\leq t/l(t)^{\delta}\ll t.$
Next, recall Lemma 4.1 with $N=2$ and error terms ${\cal R}_{1},{\cal R}_{2}$ and ${\cal R}_{3}$. From the generator calculation $\big(\rho_{1}(t)\big)^{\prime}=\sum_{x\in{\mathbb{Z}}}p(-x){\mathbb{E}}[\xi_{t}(0)\xi_{t}(x)]$ as in (96), we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\rho_{1}(t)=\rho_{1}(t-s)^{2}\sum_{x}p(x){\mathbb{P}}(\hat{\tau}_{x}>s)\hskip 1.0pt\big(1+{\cal R}_{1}+{\cal R}_{2}+{\cal R}_{3}\big).$ (106)
@@
Since the error terms have different expressions for $\alpha=1$ and $\alpha<1$, for convenience below we focus on $\alpha=1$ and $l(\infty)=\infty$, namely the recurrent case. But it is easy to check that all arguments, with slight modifications, work for $\alpha<1$. By Lemma 4.1 and the form of $s$,
${\cal R}_{1}=O\Big(\frac{1}{l(t)^{1-\theta}}\Big),\quad{\cal R}_{2}=O\Big(\frac{1}{\ell(t)^{1-\varepsilon}}\Big),\quad\mbox{and}\quad{\cal R}_{3}=O\Big(\frac{l(s)^{1-\theta}}{\ell(s)}\Big).$ (107)